OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)? A) SQUARE ROOT OF 11 B) SQUARE ROOT OF 37 C) SQUARE ROOT OF 61 D)SQUARE ROOT OF 89 PLEASE HELP ME IM TRYING REAL HARD TO GET MY MATH CREDIT.

OpenStudy (muzzack):

Place these points on a standard x-y grid or graph. Now, make a right triangle with these two points as the vertices of the hypotenuse of this triangle. Then remember "The square of the hypotenuse is equal to the sum of the squares of the other two sides" - or - c^2 = a^2 + b^2 c^2 = (-2 - 3)^2 + (1 - 7)^2 c^2 = (-5)^2 + (-6)^2 c^2 = 25 + 36 c^2 = 61 c = square root of 61 c is approximately = 7.81 <=== This is the distance between points (-2,1) and (3,7)! =========== I hope this makes sense and is helpful. or Here is how to calculate the distance between two points when you know their coordinates:

OpenStudy (muzzack):

Let us call the two points A and B

OpenStudy (anonymous):

THANK YOU SO MUCH MAN !!! YOURE THE BEST!

OpenStudy (muzzack):

We can run lines down from A, and along from B, to make a Right Angled Triangle.

OpenStudy (muzzack):

And with a little help from Pythagoras we know that: a2 + b2 = c2

OpenStudy (muzzack):

Now label the coordinates of points A and B. xA means the x-coordinate of point A yA means the y-coordinate of point A The horizontal distance "a" is (xA - xB) The vertical distance "b" is (yA - yB)

OpenStudy (muzzack):

now you get it :)

OpenStudy (anonymous):

MY TEACHER NEVER EXPLAINS ANYTHING HE DOES IT FOR ME AND THEN THE TEST COMES AND IM FREAKING LOST

OpenStudy (muzzack):

lol ill fan you and you fan me back if any help im here and welcome to openstudy

OpenStudy (anonymous):

THANKS BRO.